OpenStudy (anonymous):

Suppose we have a rectangle with one side of length 48 mm and a perimeter of 222 mm. What is the area of the rectangle?

OpenStudy (anonymous):

Let x = unknown side of rectangle. 2(48) + 2x = 222 96 + 2x = 222 2x = 126 Therefore, x = 63 Therefore area = 48(63) = 3024